what is the ratio ofionisation energy of hydrogen and be3+

what is the  ratio ofionisation energy  of hydrogen and be3+ 
 

Grade:11

1 Answers

Mahima Kanawat
5 years ago
Dear student 
Ionisation energy is directly proportional to z2/n2.
Assuming electron to be on ground state, n=1
Therefore, ratio of IE = 12/42=1/16
WITH REGARDS 
MAHIMA 
ASKIITIANS FORUM EXPERT